OpenStudy (anonymous):

One hose can fill a small swimming pool in 50 minutes. A larger hose can fill the pool in 30 minutes. How long will it take the two hoses to fill the pool working together?

OpenStudy (gorv):

1/30 +1/50=(50+30)/1500=80/1500=8/150=4/75

OpenStudy (anonymous):

18.75 min